The current in a 2.0 mm x 2.0 mm square aluminum wire is 2.5 A. What are (a) the current density and (b) the electron drift speed?
Answer #1

A)

Current Density = J = I/A

J = 2.5 / (2 * 10^-3)^2

J = 6.25 * 10^5 A/m^2

B)

V = J/nq, where v is the drift speed you are trying to find, J is the density you just found, q = the charge of an electron, and n = the number of charges carried per unit volume (the electron density).

For aluminum n = 6.02 * 10^(28) electrons/m^3
So

v = 6.25 * 10^5/ ((6.02 * 10^(28))(1.6 * 10^(-19)) in m/s

v = 0.000064887 m/s

v = 6.488 * 10^(-5) m/s